As part of preoperative care, the surgeon has ordered atropine, gr 1/300 IM. The multidose vial reads "1 mL = gr1/150." How many milliliters will be administered to the patient?
Answer #1

A s part of preoperative care, the surgeon has ordered atropine, gr 1/300 IM. The multidose vial reads "1 mL = gr1/150."

gr 1/300 = 0.2 mg

gr1/150 = 0.4 mg

The number of milliliters will be administered to the patient will be half of the vial i.e. 1ml = 0.5 ml